OpenStudy (anonymous):

[4_06]Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6.

jigglypuff314 (jigglypuff314):

were you taught point-slope form: y - y1 = m(x - x1) or slope-intercept form: y = mx + b ?

OpenStudy (anonymous):

I was taught slope intercept form

jigglypuff314 (jigglypuff314):

okay, so for slope-intercept form: y = mx + b when m =slope and b = y-intercept you were given slope = -6 so plug in -6 for m : y = -6x + b then temporarily plug in a given point to find b so if the point is (-2, -3) find b when -3 = (-6)(-2) + b then once you've found b you would plug in what you got for b back into y = -6x + b to get your equation :)

OpenStudy (anonymous):

y = -6x - 15 y = -6x - 20 y = -6x + 15 y = -6x + 20 I have to pick from these and I'm still having difficulty

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &(-2\quad ,&-3) \end{array} \\\quad \\ slope = m= -6 \\ \quad \\ y-y_1=m(x-x_1)\quad \textit{plug in your values and solve for "y"}\)

OpenStudy (texaschic101):

y = mx + b slope(m) = -6 (-2,-3) x = -2 and y = -3 now we sub y = mx + b -3 = -6(-2) + b -3 = 12 + b (subtract 12 from both sides) -3 - 12 = b -15 = b y = -6x -15